# Leveraging the true depth of LLMs

### v2

@@ 4 Efficient Parallelization of Blocks
If we naively tried to run two attention or MLP blocks by fusing their execution on a single GPU, we would not see any noticeable improvement on the inference speed. This is due to the intensive compute and memory bandwidth requirements of LLMs. For this reason we focus our attention on the tensor parallel setting, where each module’s weights are split over two or more GPUs. Rearranging the computational graph of two contiguous layers like in Fig. 2b effectively reduces the number of inter-GPU synchronizations by half. Now, each divergent path is computed in parallel over multiple accelerators, and only the single intermediate and final results need to be synchronized. While this approach is not numerically equivalent to (PAR), we nonetheless—and quite surprisingly— show that it works well on already trained models, circumventing the need to train from scratch.
LP Multi-Head Attention. Traditional tensor parallelism in MHA distributes attention heads evenly across GPUs[37], performing self-attention and output projection locally before gathering results through an all-reduce summation. Each GPU processes tensors of dimensions $Q,K,V,att\in\mathbb{R}^{T\times\frac{D}{g}}$, where $T$ is sequence length, $D$ is feature dimension, and $g$ is the number of parallel workers. The local output projection produces $o_{i}\in\mathbb{R}^{T\times D}$ for each worker $i$, with rank $\frac{D}{g}$, and then the all-reduce operation restores full rank. To implement LP, we increase the depth of the query, key, and value weight matrices ($W_{Q},W_{K},W_{V}\in\mathbb{R}^{(g_{n}\cdot h_{d})\times D}$) and widen the output projection ($W_{O}\in\mathbb{R}^{D\times(n_{h}\cdot h_{d})}$), where $n_{h}$ represents heads per GPU and $h_{d}$ is head dimensionality. The reduction operation now will simultaneously compute the full-rank output projections and the sum of all parallel layers (Fig. 5).
@@
Handling LN. Since we asume at least one GPU per parallelized layer, we can assign each original LN to the divergent path that contains the attention and FFN blocks from its original layer.
@@ 5 Experiments & Results
Figure 6: Perplexity when running pairs of consecutive layers in parallel. Perplexity of Llama2 7B and Llama3.2 3B models on the test set of RedPajama[1] when applying Layer Parallelism to $\Delta$ consecutive layers. The parallelized interval for each data point is [end indexΔ,end $[\text{end index}-\Delta,\text{end index}]$.
Table 1: 5-shot In-Context Learning accuracies across standard benchmarks. Effective Depth shows the minimum number of sequential operations from input to output. Par. Layers indicates the range of consecutive layers where pairs were processed in with Layer Parallelism. *ifeval was evaluated on 0-shot performance.
Eff. Depth MMLU PiQA Arc E. Arc C. WinoG OBQA hswag GSM8K ifeval* Avg.
Llama 2 7B
32 (Base) 0.4583 0.8009 0.8106 0.5196 0.7411 0.4520 0.7821 0.6521
27 (Ours) 0.4625 0.7933 0.8005 0.5094 0.7348 0.4600 0.7782 0.6484
26 (Ours) 0.4588 0.7927 0.7976 0.4983 0.7340 0.4460 0.7745 0.6431
25 (Ours) 0.4532 0.7851 0.7917 0.4949 0.7340 0.4440 0.7673 0.6386
24 (Ours) 0.4083 0.7845 0.7841 0.4839 0.7190 0.4360 0.7578 0.6248
23 (Ours) 0.3519 0.7829 0.7677 0.4488 0.6922 0.4240 0.7368 0.6006
Llama 3.2 3B
28 (Base) 0.5610 0.7992 0.7807 0.4872 0.7214 0.4520 0.7557 0.6510
24 (Ours) 0.5508 0.7856 0.7521 0.4753 0.7167 0.4420 0.7384 0.6373
23 (Ours) 0.5481 0.7748 0.7399 0.4735 0.7119 0.4200 0.7303 0.6284
22 (Ours) 0.4693 0.7666 0.7264 0.4497 0.6914 0.4180 0.7193 0.6058
21 (Ours) 0.3890 0.7519 0.6839 0.4061 0.6638 0.4020 0.6847 0.5690
20 (Ours) 0.3107 0.7416 0.6481 0.3652 0.6227 0.3620 0.6407 0.5273
@@
32 (Base) 0.4727 0.7769 0.7963 0.4906 0.7206 0.3300 0.5887 0.2297 0.4365 0.5380
27 (Ours) 0.4746 0.7709 0.7803 0.4881 0.7182 0.3380 0.5839 0.2221 0.4736 0.5389
26 (Ours) 0.4724 0.7666 0.7748 0.4701 0.7151 0.3400 0.5779 0.1903 0.4856 0.5325
25 (Ours) 0.4767 0.7633 0.7719 0.4650 0.7009 0.3320 0.5724 0.1478 0.4796 0.5233
24 (Ours) 0.4547 0.7655 0.7567 0.4369 0.6788 0.3040 0.5582 0.0963 0.4149 0.4962
23 (Ours) 0.4271 0.7524 0.7428 0.4113 0.6654 0.3080 0.5406 0.0697 0.3765 0.4771
@@
28 (Base) 0.5956 0.7709 0.7929 0.4650 0.7024 0.3080 0.5253 0.6475 0.6715 0.6088
24 (Ours) 0.5914 0.7622 0.7660 0.4488 0.7009 0.3000 0.5131 0.4587 0.6031 0.5716
23 (Ours) 0.5905 0.7546 0.7618 0.4471 0.6898 0.2780 0.5116 0.3571 0.5995 0.5544
22 (Ours) 0.5612 0.7530 0.7504 0.4573 0.6677 0.2880 0.5088 0.1001 0.5564 0.5159
21 (Ours) 0.5072 0.7470 0.7256 0.4019 0.6409 0.2840 0.4947 0.0311 0.4904 0.4803
20 (Ours) 0.4070 0.7410 0.7075 0.3942 0.6275 0.2900 0.4755 0.0311 0.4448 0.4576
@@
36 (Base) 0.7016 0.7644 0.8476 0.5879 0.6630 0.3720 0.5277 0.8499 0.4209 0.6372
31 (Ours) 0.6887 0.7459 0.8182 0.5367 0.6598 0.3520 0.5010 0.5375 0.3549 0.5772
30 (Ours) 0.6749 0.7497 0.8184 0.5239 0.6511 0.3340 0.4895 0.3677 0.3813 0.5545
29 (Ours) 0.6356 0.7470 0.7988 0.5043 0.6322 0.3300 0.4725 0.1797 0.3657 0.5184
28 (Ours) 0.5395 0.7432 0.7929 0.4898 0.6275 0.3000 0.4524 0.1266 0.3465 0.4909
27 (Ours) 0.4409 0.7296 0.7668 0.4454 0.5967 0.2960 0.4322 0.0356 0.2782 0.4468
26 (Ours) 0.3795 0.7062 0.7231 0.4002 0.5675 0.2840 0.4064 0.0318 0.3165 0.4239
@@
40 (Base ) 0.7883 0.8090 0.8775 0.6621 0.7451 0.4040 0.6133 0.8226 0.4652 0.6875
35 (Ours) 0.7792 0.7982 0.8695 0.6468 0.7356 0.3880 0.5825 0.6164 0.4376 0.6504
34 (Ours) 0.7742 0.7889 0.8561 0.6271 0.7388 0.3780 0.5792 0.7536 0.4700 0.6629
33 (Ours) 0.7588 0.7905 0.8561 0.6058 0.7269 0.3740 0.5693 0.6983 0.4376 0.6464
32 (Ours) 0.7370 0.7889 0.8481 0.5973 0.7103 0.3680 0.5575 0.6467 0.4017 0.6284
31 (Ours) 0.7156 0.7856 0.8552 0.5998 0.6859 0.3640 0.5412 0.4822 0.3801 0.6011
30 (Ours) 0.6543 0.7758 0.8384 0.5589 0.6732 0.3460 0.5197 0.2343 0.3657 0.5518
29 (Ours) 0.5954 0.7688 0.8215 0.5256 0.6575 0.3260 0.4945 0.1585 0.3381 0.5207
Table 2: Benchmark accuracy restoration through finetuning on Qwen3 4B with an effective depth of 27.
Finetuning Steps MMLU Arc C. GSM-8K (%)
0 (Ours) 0.4409 0.4454 0.0356
1024 (Ours) 0.6063 0.4932 0.4011
4096 (Ours) 0.6196 0.5102 0.4829
8192 (Ours) 0.6134 0.5162 0.4943
Qwen3 4B (Base) 0.7014 0.5887 0.8514
Figure 7: Wall clock time to complete the following inference tasks: KV Cache pre-filling, autoregressive generation, and single token generation with a pre-filled KV Cache. $\Delta$ indicates how many layers have been merged using LP (e.g. a $\Delta$ of 4 indicates that 2 groups of 2 layers have been converted to 2 effective layers). The gains in inference speed are directly proportional to the grade of LP. The 1-token generation task for Llama 3.2 3B does not saturate the GPU compute until a sequence length of 2048. Even in this regime, LP benefits from considerable speed-ups.
@@
Experimental protocol. For all our experiments, we use a node with two A100 SXM4 80Gb GPUs, four AMD EPYC 7742 CPUs, and 512Gb of RAM. We test for varying sequence lengths, up to $4096$ (Llama’s context window), with a batch size of $1$ unless indicated otherwise. We consider two models of the Llama family: Llama2 7B, and Llama3.2 3B, as well as two sizes from Qwen3: 4B and 14B. Given a desired effective depth, we replace the required number of normal layers with LP layers. For Llama 2 7B and 3.2 3B, the LP layers are selected based on the configuration that minimized the PPL for a given amount of LP (Fig. 6). For Qwen3, LP is applied until the 4th to last decoder layer. The rest of the layers implement the tensor parallel approach as described in [37]. For evaluation, we measure the ICL 5-shot accuracies using the lm-eval package [38]. We test the ICL accuracy of the models on several tasks: MMLU [39], PiQA [40], ARC Easy, ARC Challenge, Winogrande [41], OpenBookQA [42], Hellaswag [43], GSM-8K [44] and ifeval [45]. The perplexity of the models is always evaluated against a subset of the test set of RedPajama [1].
@@
It is worth noting that, unlike the other benchmarks, GSM-8K already drops severely in performance when applying low amounts of LP. Recent mechanistic interpretability research shows that LLMs have special circuitry for math operations, localized in a small set of parameters [46, 47]. [48] identify math specific parameters, and report a drop in accuracy of 17% when pruning them. We believe that the changes in the computational graph by the use of LP in some of the late layers of the LLM interfere with these fragile and sparse subnetworks, while leaving general language competence largely intact.
Impact on the inference speed. We run an ablation over several configurations and input sequence lengths on Figure 7 to test the speed on three different tasks: KV-Cache pre-filling, autoregressive generation up to the sequence length(with KV-Cache) and 1-token generation with a pre-filled KV-Cache of the corresponding sequence length. Our ablations show that the speed gain is directly proportional to the reduction of the effective depth of the model. For the effective depths of 25 ($\Delta=14$) in Llama 2 7B, we observe an average speed-up of 1.29x at the largest sequence length in the 1-token generation task. Likewise, for an effective depth of 23 ($\Delta=10$) in Llama 3.2 3B, we report a speed-up of 1.22x. For more aggressive parallelism, $\Delta=18$ and $\Delta=16$, we report a speed-up of 1.38x and 1.35x , at the expense of a large drop in ICL accuracy.
Fine-tuning for performance recovery. While LP provides speed improvements, associated architectural modifications may degrade model performance. To counteract this, we explored whether fine-tuning could effectively restore the original model’s capabilities. Specifically, we applied LP to layers 14–32 of Qwen3 4B, subsequently fine-tuning only these modified layers on randomly selected samples from the RedPajama training dataset [1]. Employing a batch size of 32, a linear learning rate schedule starting at $1e{-4}$, and the AdamW optimizer [49], we observed significant restoration of the model’s performance. Notably, GSM-8K accuracy improved from near-zero levels to more than half of the original model’s accuracy. It is possible that additional fine-tuning, or smarter tuning strategies, could yield further improvement, but resource constraints limited the scope of our experiments.

### v3

@@ 4 Efficient Parallelization of Blocks
Naively trying to fuse two attention or MLP sub-blocks does not result in a noticeable improvement of inference speed for large batch sizes and sequence lengths, since in these situations, inference approaches the the compute-bound regime. For this reason we focus our attention on the Tensor Parallel setting, where each module’s weights are split over two or more GPUs. Rearranging the computational graph of two contiguous layers like in Fig. 2b effectively reduces the number of inter-GPU synchronizations by half. Now, each divergent path is computed in parallel over multiple accelerators, and only the single intermediate and final results need to be synchronized. While this approach is not numerically equivalent to (PAR), we nonetheless—and quite surprisingly— show that it works well on already trained models, circumventing the need to train from scratch.
LP Multi-Head Attention. Traditional tensor parallelism in MHA distributes attention heads evenly across GPUs (shoeybi2020megatronlmtrainingmultibillionparameter), performing self-attention and output projection locally before gathering results through an all-reduce summation. Each GPU processes tensors of dimensions $Q,K,V,att\in\mathbb{R}^{T\times\frac{D}{g}}$, where $T$ is sequence length, $D$ is feature dimension, and $g$ is the number of parallel workers. The local output projection produces a low-rank $o_{i}\in\mathbb{R}^{T\times D}$ for each worker $i$, and then a final all-reduce operation performs $o=\sum_{i}^{g}o_{i}$, computing the final output.
To implement LP, we increase the depth of the query, key, and value weight matrices ($W_{Q},W_{K},W_{V}\in\mathbb{R}^{(g_{n}\cdot h_{d})\times D}$) and widen the output projection ($W_{O}\in\mathbb{R}^{D\times(n_{h}\cdot h_{d})}$), where $n_{h}$ represents heads per GPU and $h_{d}$ is head dimensionality. The reduction operation now will simultaneously compute the full-rank output projections and the sum of all parallel layers (Fig. 6). Note that this approach requires that $n_{h}$ is divisible by the total number of GPUs.
@@
Handling of the LayerNorms. Since we assume at least one GPU per parallelized layer, we can assign each original LayerNorm to the divergent path that contains the attention and FFN blocks from its original layer. We have also observed that using the same merged LayerNorm with linear interpolation and spherical linear interpolation in each divergent path yields good results. For the sake of simplicity, we conduct all of our experiments using the original LayerNorms on each divergent path.
@@ 5 Experiments & Results
[…]
Table 1: 5-shot In-Context Learning accuracies across standard benchmarks. Effective Depth shows the minimum number of sequential operations from input to output after applying LP. We use the ablation on Fig 7 to choose the LP configurations that minimized the perplexity. *ifeval was evaluated on 0-shot performance.
Eff. Depth Speed Avg Rel. MMLU PiQA Arc E. Arc C. WinoG OBQA hswag GSM8K ifeval*
@@
32 (base) x1.00 53.80 1.00 47.27 0.4 77.69 1.0 79.63 0.8 49.06 1.5 72.06 1.3 33.00 2.1 58.87 0.5 22.97 1.2 43.65
27 (ours) x1.15 53.89 1.00 47.46 0.4 77.09 1.0 78.03 0.9 48.81 1.5 71.82 1.3 33.80 2.1 58.39 0.5 22.21 1.1 47.36
26 (ours) x1.19 53.25 0.99 47.24 0.4 76.66 1.0 77.48 0.9 47.01 1.5 71.51 1.3 34.00 2.1 57.79 0.5 19.03 1.1 48.56
25 (ours) x1.23 52.33 0.97 47.67 0.4 76.33 1.0 77.19 0.9 46.50 1.5 70.09 1.3 33.20 2.1 57.24 0.5 14.78 1.0 47.96
24 (ours) x1.28 49.62 0.92 45.47 0.4 76.55 1.0 75.67 0.9 43.69 1.5 67.88 1.3 30.40 2.1 55.82 0.5 9.63 0.8 41.49
23 (ours) x1.31 47.71 0.89 42.71 0.4 75.24 1.0 74.28 0.9 41.13 1.4 66.54 1.3 30.80 2.1 54.06 0.5 6.97 0.7 37.65
@@
28 (base) x1.00 60.88 1.00 59.56 0.4 77.09 1.0 79.29 0.8 46.50 1.5 70.24 1.3 30.80 2.1 52.53 0.5 64.75 1.3 67.15
24 (ours) x1.12 57.16 0.94 59.14 0.4 76.22 1.0 76.60 0.9 44.88 1.5 70.09 1.3 30.0 2.1 51.31 0.5 45.87 1.4 60.31
23 (ours) x1.15 55.44 0.91 59.05 0.4 75.46 1.0 76.18 0.9 44.71 1.5 68.98 1.3 27.80 2.0 51.16 0.5 35.71 1.3 59.95
22 (ours) x1.19 51.59 0.85 56.12 0.4 75.30 1.0 75.04 0.9 45.73 1.5 66.77 1.3 28.80 2.0 50.88 0.5 10.01 0.8 55.64
21 (ours) x1.23 48.03 0.79 50.72 0.4 74.70 1.0 72.56 0.9 40.19 1.4 64.09 1.4 28.40 2.0 49.47 0.5 3.11 0.5 49.04
20 (ours) x1.28 45.76 0.75 40.70 0.4 74.10 1.0 70.75 0.9 39.42 1.4 62.75 1.4 29.00 2.0 47.55 0.5 3.11 0.5 44.48
@@
36 (base) x1.00 63.72 1.00 70.16 0.4 76.44 1.0 84.76 0.8 58.79 1.5 66.30 1.3 37.20 2.0 52.77 0.5 84.99 1.0 42.09
31 (ours) x1.13 57.72 0.91 68.87 0.4 74.59 1.0 81.82 0.8 53.67 1.5 65.98 1.3 35.20 2.2 50.10 0.5 53.75 1.0 35.49
30 (ours) x1.15 55.45 0.87 67.49 0.4 74.97 1.0 81.84 0.9 52.39 1.4 65.11 1.3 33.40 2.1 48.95 0.5 36.77 0.9 38.13
29 (ours) x1.18 51.84 0.81 63.56 0.4 74.70 1.0 79.88 0.9 50.43 1.4 63.22 1.4 33.00 2.2 47.25 0.5 17.97 0.8 36.57
28 (ours) x1.21 49.09 0.77 53.95 0.4 74.32 1.0 79.29 0.9 48.98 1.6 62.75 1.3 30.0 2.1 45.24 0.5 12.66 0.5 34.65
27 (ours) x1.25 44.68 0.70 44.09 0.4 72.96 1.0 76.68 0.9 44.54 1.5 59.67 1.4 29.60 2.0 43.22 0.5 3.56 0.5 27.82
@@
40 (base) x1.00 68.75 1.00 78.83 0.4 80.90 1.0 87.75 0.8 66.21 1.4 74.51 1.3 40.40 2.1 61.33 0.5 82.26 1.2 46.52
35 (ours) x1.12 65.04 0.95 77.92 0.4 79.82 1.0 86.95 0.9 64.68 1.5 73.56 1.3 38.80 2.0 58.25 0.5 61.64 1.1 43.76
34 (ours) x1.15 66.29 0.96 77.42 0.4 78.89 1.0 85.61 0.9 62.71 1.5 73.88 1.3 37.80 2.0 57.92 0.5 75.36 1.0 47.00
33 (ours) x1.18 64.64 0.94 75.88 0.4 79.05 1.0 85.61 0.9 60.58 1.4 72.69 1.4 37.40 2.2 56.93 0.5 69.83 1.0 43.76
32 (ours) x1.21 62.84 0.91 73.70 0.4 78.89 1.0 84.81 0.9 59.73 1.4 71.03 1.4 36.80 2.0 55.75 0.5 64.67 0.9 40.17
31 (ours) x1.24 60.11 0.87 71.56 0.4 78.56 1.0 85.52 0.9 59.98 1.5 68.59 1.3 36.40 2.0 54.12 0.5 48.22 0.8 38.01
30 (ours) x1.27 55.18 0.80 65.43 0.4 77.58 1.0 83.84 0.9 55.89 1.5 67.32 1.3 34.60 2.1 51.97 0.5 23.43 0.8 36.57
Table 2: Benchmark accuracy restoration for LP-applied Qwen3 models. Fine-tuned entries use 4096 additional training steps; left columns report the fine-tuned accuracy, and right columns report the no-finetune/baseline values.
Model Eff. depth Fine-tuned No fine-tuning
MMLU Arc C. GSM-8K (%) MMLU Arc C. GSM-8K (%)
Qwen3 4B 36 (Baseline) 70.16 0.4 58.79 1.5 84.99 1.0
31 69.11 0.4 54.10 1.5 62.47 1.3 68.87 0.4 53.67 1.5 53.75 1.0
30 68.43 0.4 54.10 1.5 56.56 1.4 67.49 0.4 50.43 1.4 36.77 0.9
27 61.96 0.4 51.02 1.6 48.29 1.4 44.09 0.4 44.54 1.5 3.56 0.5
Qwen3 14B 40 (Baseline) 78.83 0.4 66.21 1.4 82.26 1.2
35 77.89 0.3 63.05 1.4 81.73 1.1 77.92 0.4 64.68 1.5 61.64 1.1
32 74.38 0.4 59.30 1.4 71.27 1.3 73.70 0.3 59.73 1.4 64.67 0.9
Figure 8: Wall clock time to complete the following inference tasks: KV Cache pre-filling, autoregressive generation, and single token generation with a pre-filled KV Cache. $\Delta$ indicates how many layers have been merged using LP (e.g. a $\Delta$ of 4 indicates that 2 groups of 2 layers have been converted to 2 effective layers). The gains in inference speed are roughly proportional to the amount of LP. The 1-token generation task for Llama 3.2 3B does not saturate the GPU compute until a sequence length of 2048. Even in this regime, LP benefits from considerable speed-ups.
@@
Experimental protocol. For all our experiments, we use a node with x2 A100 SXM4 80Gb GPUs, x4 AMD EPYC 7742 CPUs, and 512Gb of RAM. We test for varying sequence lengths, up to $4096$ (Llama’s context window), with a batch size of $1$ unless indicated otherwise. We consider two models of the Llama family: Llama2 7B, and Llama3.2 3B, as well as two sizes from Qwen3: 4B and 14B. Given a desired effective depth, we replace the required number of normal layers with LP layers. For Llama 2 7B and 3.2 3B, the LP layers are selected based on the configuration that minimized the PPL for a given amount of LP (Fig. 7). For Qwen3, LP is applied until the 4th to last decoder layer. The rest of the layers implement the tensor parallel approach as described in (shoeybi2020megatronlmtrainingmultibillionparameter). For evaluation, we measure the ICL 5-shot accuracies using the lm-eval package (eval-harness). We test the ICL accuracy of the models on several tasks: MMLU (hendrycks2021measuringmassivemultitasklanguage), PiQA (bisk2019piqareasoningphysicalcommonsense), ARC Easy, ARC Challenge, Winogrande (sakaguchi2021winogrande), OpenBookQA (mihaylov2018suitarmorconductelectricity), Hellaswag (zellers2019hellaswag), GSM-8K (cobbe2021trainingverifierssolvemath) and ifeval (zhou2023instructionfollowingevaluationlargelanguage). The perplexity of the models is always evaluated against a subset of the test set of RedPajama (together2023redpajama).
@@
It is worth noting that, unlike the other benchmarks, GSM-8K already drops severely in performance when applying low amounts of LP. Recent mechanistic interpretability research shows that LLMs have special circuitry for math operations, localized in a small set of parameters (stolfo2023mechanistic; yu2024interpreting). (christ2025mathneurosurgeryisolatinglanguage) identify math specific parameters, and report a drop in accuracy of 17% when pruning them. We hypothesize that the changes in the computational graph by the use of LP in some of the late layers of the LLM interfere with these fragile and sparse subnetworks, while leaving general language competence largely intact.
Impact on the inference speed. We run an ablation over several configurations and input sequence lengths on Figure 8 to test the speed on three different tasks: KV-Cache pre-filling, autoregressive generation up to the sequence length(with KV-Cache) and 1-token generation with a pre-filled KV-Cache of the corresponding sequence length. Our ablations show that the speed gain is strongly correlated with the reduction of the effective depth of the model. For the effective depths of 25 ($\Delta=14$) in Llama 2 7B, we observe an average speed-up of 1.29x at the largest sequence length in the 1-token generation task. Likewise, for an effective depth of 23 ($\Delta=10$) in Llama 3.2 3B, we report a speed-up of 1.22x. For more aggressive parallelism, $\Delta=18$ and $\Delta=16$, we report a speed-up of 1.38x and 1.35x , at the expense of a large drop in ICL accuracy.
Fine-tuning for performance recovery. While LP provides speed improvements, associated architectural modifications may degrade model performance. To counteract this, we explored whether fine-tuning could effectively restore the original model’s capabilities. We apply LP to some configurations of Qwen3-4B and Qwen3-14B (Table 2), and fine-tune the LP layers on randomly selected samples from the RedPajama training set (together2023redpajama). We employ a batch size of 32, a linear learning rate schedule starting at 1e41e{-4} and the AdamW optimizer (loshchilov2017decoupled). We observe a significant restoration of the benchmark accuracies for Qwen3-4B with an effective depth of 27, especially on GSM-8K, which recovered from near-zero levels. Less aggressive usage of LP results in a less pronounced recovery of the accuracy, and fails to fully recover the original model’s performance. It is possible that additional fine-tuning, or smarter tuning strategies, could yield further improvements, but resource constraints limited the scope of our experiments.